Two coils of resistance 3 Ohm and 6 Ohm . are connected in series across a battery of emf 12v . find electrical energy consumed in 1 min in each resistance when these are connected in series?

Grade:10

1 Answers

Arun
5 years ago
we know that 
power= energy/time 
so 
we will find the power first 
we know that power = V²/R
so now we will find the Reff.
as they are connected in series then effective resistance= 3+6 = 9ohm.
so now power = 12²/9 = 16.
so power = work done(energy) / time 
hence energy= 16× 60. = 960jule
 
 
Regards
Arun